# Scalable Private Search with Wally

### v5

@@ 3 System Overview
Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “washes” queries of identifying information.
Figure 1 shows a high-level design of Wally. It consists of three entities: the server, the clients, and the anonymization network. Wally runs in epochs of time; within each epoch, a fixed number of clients make queries. Within each epoch, Wally performs a search that outputs the most relevant results to client queries, similar to insecure search engines. Wally does not require any synchronization across clients. Each client can independently decide to participate or skip an epoch. However, we require a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. Overall, the privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated. Wally’s privacy notion also requires that each client makes at most $M$ queries in the system’s lifetime; therefore, a client can participate at most $M$ epochs. The correctness guarantee is that all the queries within epoch are processed by the end of the epoch.
In Wally each epoch is divided into time slots, each a second long, e.g., and privacy is guaranteed by the following:
  • Hiding the content of the queries (somewhat homomorphic encryption).
  • Hiding the real queries of clients among fake queries (differential privacy).
  • Removing the identifiable information from each query and shuffling the batch of queries received within a time slot (anonymization network).
  • Hiding the information revealed by queries order and time of arrival (random query schedule).
Embedding-based search.
@@
Similar to state of the art insecure semantic search and previous secure search, Wally utilizes (vector) embeddings for semantic search. Embeddings are a machine learning-based technique that map unstructured objects like documents, images, and videos to $d$-dimensional vectors of floats. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, they are used to find relevant entries in search systems. Embeddings are usually much smaller than the objects themselves, which results in a significant reduction in client-to-server communication. Embeddings can be generated for different data types, including text, image, and video. Therefore, Wally can support all these data types. Wally is compatible with all the recent embedding models, such as BERT [12].
In the following sections, we introduce the insecure version of semantic search in Wally. We then highlight the privacy issues in this version. Finally, we discuss how we overcome these privacy issues.
3.1 Insecure Search Pipeline
@@
In the offline phase, the server maps each document to the corresponding embedding. At the time of a query the client embeds the query into an embedding. Once the client query and the server documents are represented as embeddings, Wally uses an approximate nearest neighbor algorithm to find the documents most relevant to the client’s query. Our two-step approximate nearest neighbor protocol has the following flow:
@@
  • Server initialization. At initialization, the server divides the documents (using their embeddings) into $K$ clusters using K-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$. The server sends the centroid embeddings to the client.
@@
  • Pick nearest centroid. To make a query, the client first locally finds the cluster nearest to the query embedding. For this, the client computes a similarity between the query and each of the centroid embeddings and picks a cluster whose centroid has high similarity.
  • Server computation. The client then sends the nearest cluster index $C^{\prime}$ and query embedding $q^{\prime}$ to the server. The server computes similarity between the query embedding and each embedding in cluster $C^{\prime}$. The server returns the similarity scores to the client.
  • Pick nearest entry. The client sorts these scores and picks the entry closest to the query.
  • Metadata retrieval. The client then fetches the metadata associated with the closest entry from the server.
@@
The clustering technique provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is in a different cluster than the cluster with the closest centroid. Therefore, finding a balance is important for the overall usability of the system.
The above protocol is not private. To ensure that the protocol does not reveal any information about the client’s query, it must hide the following information from the server:
@@
  • Nearest cluster. The client reveals to the server the nearest cluster ID. Recall that clustering puts highly correlated documents together in a cluster. Therefore, the cluster ID can reveal the kind of data the client holds. For example, if the client query is a song embedding, the cluster ID might reveal the genre of the song.
  • Closest entry index. To fetch the relevant metadata, the client sends the index of the closest entry to the server. Learning this could reveal server information about the query.
We show how we hide the above data below and in Section 6.
3.2 Hiding the Query Embedding
@@
We rely on efficient vectorized, RLWE-based SHE described in Section 2 to hide a query embedding. At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster. The server then returns encrypted scores to the client. SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient. Overall, in our implementation, the client query and server response are a single RLWE ciphertext of sizes 226kB, and 53kB, respectively, (the former with evaluation keys) and it only takes 19 ms for the server to compute similarity scores for a single cluster homomorphically.
3.3 Hiding the nearest cluster
By using SHE, the content of the query is hidden from the server, but the server learns the nearest cluster. Hiding the nearest cluster is a significant performance bottleneck in previous schemes like Tiptoe. Specifically, to conceal the nearest cluster for a single query, Tiptoe requires the server to perform SHE operations against the entire database (all the clusters). Otherwise, skipping clusters will reveal to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster in the most robust possible sense, but this results in a prohibitively high server computation cost.
Instead, we propose an approach based on differential privacy. The key advantage is that, in expectation, the server computes encrypted similarity scores only against a few clusters per query. Our approach results in a significant improvement in server computation. An additional advantage of our approach is a reduction in request size; concretely, the request size is only $O(d)$, around $\sqrt{N}$ smaller than Tiptoe. Conversely, differential privacy is a weaker privacy notion than the full, cryptographic obliviousness Tiptoe achieves.
The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making the queries at any given time. Therefore, we propose a novel mechanism to hide a particular client query among the crowd of queries from all the clients. Our proposed approach provides $(\epsilon,\delta)$-differential privacy. This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $\epsilon$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
As the first step, we require all the traffic (queries and responses) to and from the server to route through the anonymization network. Intuitively, one would expect that doing this would hide the query in the crowd of queries, as queries do not carry identifiable information. However, note that only using an anonymization network is not sufficient because by observing traffic patterns, the server can learn information about a client. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster then by observing that the cluster is accessed, the server can infer whether a request is from the client or not. Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client. For example, the queries towards each epoch’s start could belong to a particular user.
To achieve the desired privacy, we make the following two additional changes, highlighted in Figure 3:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\mathbb{E}[\mathcal{D}]$ fake queries. For each fake query, the client creates an SHE encryption of zero and pseudorandomly picks one of the clusters.
Intuitively, these two changes suppress the leakage, as mentioned above. In Section 6, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees differential privacy.
3.4 Hiding Metadata Index
After learning the index of the nearest entry, the client retrieves metadata associated with it. When metadata per entry is small, the server returns metadata of the entire cluster with each query. Note that this leaks no extra information because the server already learns the cluster ID for each query. The client then locally selects the metadata associated with the most relevant entry. However, if the metadata is quite large, the client can run Keyword PIR to fetch the metadata of only the most relevant entry within a cluster.
3.5 Keyword Search
Wally also supports keyword searches in which the client is interested in retrieving data associated with a keyword. Using a random hash function, the server divides the database into $K^{\prime}$ random buckets (instead of clusters). For each query, the client uses the same hash function to find the bucket in which the keyword falls locally and then uses the efficient keyword PIR scheme to fetch a specific record within a bucket. The differential privacy part remains the same as in semantic search. That is, clients make fake queries for random buckets at a random schedule.
4 Efficient SHE-Based Instantiation
In Wally, for semantic search, SHE is used to calculate similarity scores between the query and nearest cluster embeddings, and for the keyword search, an SHE based Keyword PIR is used to fetch the entry associated with query keyword among the entries in the bucket in which query keyword falls. In this section, we discuss the details of our efficient SHE-based instantiation.
4.1 SHE details in semantic search
The server database is encoded as an array of plaintext vectors. The output is a single response ciphertext with at most $n$ encrypted similarity scores, where $n$ is the RLWE polynomial dimension.
@@ 4.1 SHE details in semantic search
SHE dot-product computation.
Recall, from Section 2, to calculate cosine similarity between two embeddings, the server needs to compute a dot-product between them. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $n=d=4$.
@@
Recall from Section 2 that homomorphic rotation is costly. In the above explanation, the server performs $d$ rotations even after packing to align the query. In Wally  we use the standard baby-step giant-step (BSGS) optimization [23] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step). See Algorithm 1 in [23] for the BSGS algorithm.
Float-to-integer encoding.
Recall from Section 2 that vectorized SHE supports operations over vectors of integers while embeddings are vectors of floats. In Wally we convert each float into an integer by multiplying it with the appropriate scaling factor having the plaintext modulus be large enough to perform the computation over the integers. The choice of scaling factor depends on the size of SHE plaintext modulus $t$ and the required precision.
@@
4.2 SHE details in keyword search
The server database consists of $N$ (keyword, value) entries. Recall that the server divides these entries into $B$ buckets using a random hash function. Within each bucket, in Wally  we use a keyword PIR construction based on cuckoo hashing and index PIR.
At a high level, for each bucket independently, the server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at bucket index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these indices and locally find the index containing the entry and associated data. Wally uses MulPIR [4] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
Each server bucket consisting of $N/B$ entries is represented as a $\sqrt{N/B}\times\sqrt{N/B}$ matrix $\mathbf{D}$. MulPIR uses a polynomial version of SHE, where each plaintext is a polynomial. For this explanation, we assume that each entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.

### v6

@@ III Private search with inefficient communication
Figure 1: High-level overview of Wally. Throughout an epoch, clients send encrypted queries, real and fake, at random time slots to an anonymization network which “removes” identifying information.
In this section we provide a warm-up protocol that provides the required privacy but has high communication overhead. This protocol act as stepping stone towards our final protocol discussed in Section V.
Embedding.
@@
Similar to state of the art insecure semantic search and previous secure search, This protocol utilizes embeddings for search. Embeddings are a machine learning-based technique that map unstructured objects like text, images, and videos to $d$-dimensional vectors of floating-point numbers. These embeddings retain the semantic relationship between objects by placing similar objects close together in a vector space. Hence, they are used to find relevant entries in search systems. Embeddings can be generated for different data types, including text, image, and video. Therefore, our protocol can support all these data types. Wally is compatible with recent embedding models, such as BERT [30].
Clustering based nearest neighbors search.
@@
The server has a database of $N$ entries, where each entry consists of a document such as text or image and a bit string of associated metadata. In the offline phase, the server maps each document to the corresponding embedding. At the time of a query, the client embeds the query into an embedding. Once the client’s query and the server’s documents are represented as embeddings, the client’s objective is to identify the nearest server embedding and its associated metadata. Finding the exact nearest neighbor is computationally expensive even in an insecure environment. Therefore, we utilize an approximate nearest neighbor algorithm using clustering to find the most relevant documents to the client’s query. The protocol has the following flow:
@@
  • Initialization. At initialization, the server divides the document embeddings into $K$ clusters using $K$-means clustering, a heuristic-based technique that partitions data into $K$ disjoint clusters $C_{1},\cdots,C_{K}$, and outputs these clusters and their centroid embeddings $c_{1},\cdots,c_{K}$. The server also divides the document metadata into the same clusters. The server sends the centroid embeddings to the client.
@@
  • Query. To make a query, the client first locally finds $\Delta$ clusters nearest to the query embedding $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$. For this, the client computes a cosine similarity between the query and each of the centroid embeddings and picks clusters whose centroid has the highest similarity.
  • Response. The client then sends the nearest clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the server. The server responds by sending the embeddings and their metadata in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{\Delta}$ to the client.
  • Local computation. The client locally computes cosine similarity between the query embedding and each embedding in clusters $C^{\prime}_{1},\cdots,C^{\prime}_{1}={\Delta}$, sorts the scores to find the closest embedding, and the corresponding metadata.
@@
The clustering provides a trade-off between performance and accuracy; generally, increasing the number of clusters improves the performance (reduces server computation and response size). However, it degrades the accuracy because there is a higher chance that a potential nearest neighbor is missed. Therefore, finding a balance is important for the overall usability of the system.
This protocol is not private. To prevent the protocol from disclosing any information about the client’s query, it must conceal the closest clusters from the server. Remember that clustering groups highly correlated documents together in a cluster. Consequently, knowing the nearest clusters could reveal the information about the client’s query.
III-A Hiding nearest clusters using differential privacy
Hiding the nearest clusters is a significant performance bottleneck in previous schemes like Tiptoe. For instance, to conceal a single nearest cluster, Tiptoe requires the client to send an encrypted query for each cluster; similarly, it requires the server to perform SHE operations against all the clusters. This is because omitting clusters will indicate to the server that the client is not interested in them. Computing over the entire database hides the nearest cluster completely, but this results in a prohibitively high communication and computation.
We propose an approach based on differential privacy. The key advantage is that touching all the clusters is not required, which results in a significant improvement in communication and server computation than Tiptoe. We emphasize that differential privacy is a different privacy notion than the full obliviousness Tiptoe achieves; however, it is still a standard privacy notion. Intuitively, our proposed approach provides $(\epsilon,\delta)$-differential privacy. This approach guarantees to every client that the maximum privacy loss due to its queries is essentially bounded by $e^{\epsilon}$ with all but $\delta$ probability. In other words, a client’s queries do not significantly influence the outcome of any statistical inference run by the server.
The starting point of our solution is an observation that in large-scale search engines, sufficiently many clients are making queries at any given time. Therefore, we can hide a particular client query among the batch of queries from all the clients. Our approach require following steps.
Epochs.
As the first, we require the system to run in epochs of time, each one second long; within each epoch, a fixed number of clients make queries. This does not require any synchronization across clients. Each client can independently decide to participate or skip an epoch. However, we assume a minimum number of honest clients stay online throughout an epoch and that a participating client can only make at most $\Delta$ queries within each epoch. We also require that the client send $\Delta$ clusters as independent queries to the server instead of sending them together. The worst privacy loss for a particular client is the sum of the loss in each epoch in which the client has participated. The correctness guarantee is that all the queries within an epoch are processed by the end of the epoch.
Anonymized queries.
We also require all the queries to be anonymized. For this, each client will send the queries via an anonymization network.
Note that even though queries are anonymized, by observing traffic patterns, the server can still learn information about the queries. For example, if the server has side information that a particular client is interested in a specific cluster and all the other clients are known not to access the cluster, then by observing that the cluster is accessed, the server can infer whether a request is from the client or not. Therefore, we require deniability of queries. Similarly, the server can exploit the arrival times of queries to infer if queries belong to a particular client. For example, the queries towards each epoch’s start could belong to a particular user. Therefore, we need to suppress leakage through the timing channel.
To achieve the desired properties, we make the following two additional changes, highlighted in Figure 3:
  • Fake queries. In addition to real queries, each client makes a few fake queries. Specifically, for each cluster, the client makes $Z\leftarrow\mathcal{D}$ fake queries, where distribution $\mathcal{D}$ is chosen to guarantee differential privacy after fake queries from many clients are aggregated. In expectation, the client makes $K\leftarrow\mathbb{E}[\mathcal{D}]$ fake queries..
Intuitively, these two changes suppress the leakage, as mentioned above. In Section VI, we show that using the negative binomial distribution for fake queries $\mathcal{D}$ along with random query schedule guarantees $(\epsilon,\delta)$ differential privacy.
Drawback of warm-up protocol.
Even though the protocol conceals the closest clusters, it incurs high communication costs. Concretely, the server transmits embeddings and metadata for all the nearest clusters back to the client, leading to substantial communication overhead.
IV Optimizing the communication using SHE
We use SHE to reduce the communication of our warm-up protocol. In this section, we discuss the details of our efficient SHE-based instantiation.
IV-A Reducing communication in nearest neighbor
@@
We rely on vectorized, RLWE-based SHE described in Section II to hide a query embedding. At a high level, the client sends an SHE encryption of its embedding to the server, and the server homomorphically (under encryption) computes similarity scores between the query embedding and all the embeddings in the nearest cluster. The server then returns encrypted scores to the client. SHE computation is inherently expensive; we utilize various techniques to make similarity computation efficient.
@@
Secure dot-product computation.
From Section II, to calculate cosine similarity between two normalized embeddings, the server needs to compute a dot-product between them. To achieve that, the server must align the query to ensure that elements of a query vector only get multiplied with corresponding element of plaintext vectors, i.e., every element-wise multiplication should be of the form $e_{j}^{i}*q^{i}$. The server will use homomorphic rotation to move the elements within the encrypted query vector to the correct alignment. Specifically, for a given slice, starting with the first plaintext vector $j=0$, the server performs vectorized ciphertext-plaintext multiplication between the query ciphertext and the first plaintext vector $j=0$. This results in encrypted $p_{0}*\hat{q}=[e_{0}^{0}*q^{0},e_{1}^{1}*q^{1},e_{2}^{2}*q^{2},\cdots,e_{d-1}^{d-1}*q^{d-1}]$. No rotation is required for first multiplication because the query is already aligned with the first vector. The elements of the second plaintext vector $j=1$ are $p_{1}=[e_{0}^{1},e_{1}^{2},e_{2}^{3},\cdots,e_{d-1}^{0}]$. Hence, the server cannot directly multiply it with the query vector. The server first homomorphically rotates the query vector one slot to the left, resulting in query vector $\hat{q}^{1}=[q^{1},q^{2},\cdots,q^{0}]$ then multiplies the result: $p_{1}*\hat{q}^{1}=[e_{0}^{1}*q^{1},e_{1}^{2}*q^{2},e_{2}^{3}*q^{3},\cdots,e_{d-1}^{0}*q^{0}]$. The server then repeats this for all the remaining plaintext vectors in a slice: rotating the query vector in increments of one and multiplying it with the plaintext vector. By the end the server has $p_{0}*\hat{q},p_{1}*\hat{q}^{1},\cdots,p_{d-1}*\hat{q}^{d-1}$ encrypted multiplications. The server then sums all of these vectors together, resulting in a single encrypted vector in which each element is a dot product between the client query vector and one of the embeddings in the slice. Figure 2 shows an example server computation for a single slice for $n=d=4$.
@@
Recall from Section II that homomorphic rotation is costly. In the above explanation, the server performs $d$ rotations even after packing to align the query. In Wally  we use the standard baby-step giant-step (BSGS) optimization [32] to minimize the number of rotations: for $d$-dimensional inner products, BSGS requires only $\sqrt{d}$ rotations for step sizes of $1$ (baby step) and $\sqrt{d}$ (giant step). See Algorithm 1 in [32] for the BSGS algorithm.
@@
IV-B Metadata Fetch
After determining the index of the nearest entry, the client retrieves metadata related to it. When the metadata per entry is small, the server responds by returning the metadata for the entire cluster with each query. It’s important to note that this doesn’t leak any additional information since the server already knows the query cluster. The client then locally selects the metadata associated with the most relevant entry. However, if the metadata is large (in kilobytes), it becomes infeasible to download the entire cluster. In this case, the client obliviously retrieves the metadata of only the most relevant entry within a single cluster. To achieve this, we use Keyword private information retrieval (PIR) scheme based on cuckoo hashing and index PIR.
Concretely, after getting encrypted scores, the client finds the most relevant embedding locally. Use the identity of this entry to generate a Keyword PIR query. The metadata query is then clustered in which this embedding falls, as well as the Keyword PIR query.
Protecting privacy leakage.
Note that, it is still required to protect the query cluster. Therefore, clients will require participating in an additional epoch with $\Delta=1$. For this epoch also, we required queries to be anonymized, clients generate fake queries, and execute queries at random slots within an epoch.
Details of Keyword PIR.
The server encode each entry into Keyword-value pair. Where keyword equals entry ID and value is the metadata. For ease of explanation we assume that each cluster on average consist of $N^{\prime}$ entries. The server uses the cuckoo hashing to map the $N^{\prime}$ entries into a table of size $O(N^{\prime})$ using two random hash functions. The guarantee is that if a keyword is present in the database, it must be at entry index $h_{1}(\textit{keyword})$ or $h_{2}(\textit{keyword})$ in the table. The client then uses index PIR to privately fetch entries at these two indices and locally find the index containing the entry and associated data. Wally uses MulPIR [33] as the underlying index PIR, an efficient single-server PIR scheme based on SHE. We emphasize that even though we initialize Wally with MulPIR, it is compatible with any recent efficient PIR scheme.
Details of MulPIR.
The server represent cuckoo table into $\sqrt{N^{\prime}}\times\sqrt{N^{\prime}}$ matrix For this explanation, we assume that each metadata entry has the same size as SHE plaintext. If entries are larger than a plaintext, then we split each entry across multiple plaintexts, and if entries are smaller, then we pack multiple entries into each plaintext polynomial when applicable. A client query in MulPIR is a single ciphertext $\hat{q}$ encrypting row and column indices. Given a query, the server does the following.